### v3

# Detecting Out-of-Distribution Objects through Class-Conditioned Inpainting

Abstract
Recent object detectors have achieved impressive accuracy in identifying objects seen during training. However, real-world deployment often introduces novel and unexpected objects, referred to as out-of-distribution (OOD) objects, posing significant challenges to model trustworthiness. Modern object detectors are typically overconfident, making it unreliable to use their predictions alone for OOD detection. To address this, we propose leveraging an auxiliary model as a complementary solution. Specifically, we utilize an off-the-shelf text-to-image generative model, such as Stable Diffusion, which is trained with objective functions distinct from those of discriminative object detectors. We hypothesize thatbu this fundamental difference enables the detection of OOD objects by measuring inconsistencies between the models. Concretely, for a given detected object bounding box and its predicted in-distribution class label, we perform class-conditioned inpainting on the image with the object removed. If the object is OOD, the inpainted image is likely to deviate significantly from the original, making the reconstruction error a robust indicator of OOD status. Extensive experiments demonstrate that our approach consistently surpasses existing zero-shot and non-zero-shot OOD detection methods, establishing a robust framework for enhancing object detection systems in dynamic environments.
Machine Learning, ICML
1 Introduction
Figure 1: Intuition behind RONIN for OOD detection. Object detectors can overconfidently make wrong predictions on unseen objects based on a pre-defined set of ID labels; in this case, predicting two OOD deers as “sheep”. RONIN leverages the label predictions to condition the resynthesizing process of an off-the-shelf text-to-image diffusion model, producing similar inpaintings for correct predictions and dissimilar inpaintings for incorrect ones. Under a similarity measurement, RONIN can identify the wrong prediction of unseen objects, therefore ODD detection.
Object detection systems have achieved remarkable progress in recent years, becoming integral to a wide range of applications in both online and offline settings. These include, but are not limited to, environmental monitoring (Beery et al., 2019), industrial manufacturing (Ahmad & Rahimi, 2022), and healthcare (Ragab et al., 2024). As these systems increasingly influence decision-making processes, ensuring their reliability and robustness has become a key challenge.
One common failure mode arises when object detectors encounter object categories not seen during training, referred to as out-of-distribution (OOD) cases. Such occurrences are inevitable due to the dynamic and ever-changing nature of the real world. Unfortunately, modern object detectors often exhibit overconfidence, frequently misclassifying OOD objects as belonging to in-distribution (ID) categories, thereby resulting in false positives (Pathiraja et al., 2023). This overconfidence underscores a fundamental limitation: the internal responses of object detectors cannot reliably identify these errors. Consequently, alternative strategies are needed to enhance the trustworthiness of these systems.
A popular strategy is to “rehearse” these error cases during detector training. For example, Munir et al. (2022); Pathiraja et al. (2023) introduced auxiliary losses to regularize model confidence, while Du et al. (2022b, a) proposed methods to learn more distinguishable detector features between ID and OOD categories. While promising, these strategies require retraining and are not widely implemented in pre-trained, publicly accessible object detectors. Given that many stakeholders rely on off-the-shelf detectors for their applications, there is a strong demand for effective post-hoc solutions to address OOD detection challenges.
In this paper, we propose leveraging an additional pre-trained model to provide auxiliary information for effective OOD detection. Specifically, we utilize off-the-shelf text-conditioned generative models (Rombach et al., 2021; Saharia et al., 2022), which are trained to synthesize realistic images from text prompts, such as category names. We hypothesize that the distinct training objectives of object detectors and generative models lead to inconsistent responses for OOD cases. These inconsistencies thus serve as a robust post-hoc indicator for OOD detection, offering a practical and accessible solution to enhance the reliability of object detection systems without requiring retraining.
More specifically, given a detected object bounding box and its predicted ID class label by the object detector, we propose performing class-conditioned inpainting for OOD detection. We mask the image portion within the box and use the ID class label as the prompt to resynthesize it. If the masked-out object is ID, the reconstructed image portion should be similar to the original, both visually and semantically. In contrast, for an OOD object, the inpainted portion prompted by the misclassified ID label will significantly deviate from the original. This discrepancy provides a valuable cue for distinguishing ID and OOD objects, requiring neither retraining nor access to ID or OOD training examples. We name our approach ZeRo-shot CONditional INpainting (RONIN), as illustrated in Figure 1.
We evaluate RONIN on real-world benchmark datasets, including PASCAL VOC (Everingham et al., 2010), BDD-100K (Yu et al., 2020), MS-COCO (Lin et al., 2014), and OpenImages (Kuznetsova et al., 2020), covering a broad range of categorical domains and context scenarios, from indoor scenes to in-the-wild environments, for comprehensive evaluation of our proposed method. The empirical results demonstrate RONIN’s superior OOD detection performance compared to existing methods. Notably, RONIN excels in identifying near-OOD cases, which have semantic meanings closely related to some ID categories and are therefore particularly challenging to detect. This improvement is achieved by incorporating potential near-OOD labels as exclusions, resulting in larger reconstruction errors when the original object belongs to one of these near-OOD labels. Extensive ablation studies, analyses, and visualizations further validate the effectiveness of RONIN.
In summary, RONIN offers three notable advantages, positioning itself as a plug-and-play solution for OOD detection:
  • Retraining-free and zero-shot: RONIN requires no access to training data for either ID or OOD categories, making it highly versatile and data-efficient.
  • Ease of implementation and compatibility: RONIN is straightforward to implement and compatible with most existing object detectors, including open-vocabulary object detectors (Liu et al., 2023a), which still rely on a pre-defined set of ID concepts to limit their scope.
  • Leveraging off-the-shelf generative models: By utilizing pre-trained generative models, RONIN continuously benefits from advancements in generative modeling, ensuring improved effectiveness and efficiency over time.
Remark.   While straightforward to implement, RONIN is far from a trivial contribution. It leverages the inconsistency between discriminative and generative models as an effective post-hoc cue for identifying OOD objects. Compared to Liu et al. (2023b), RONIN extends its applicability to object detection and eliminates the need to train a diffusion model on ID data. With the growing accessibility of pre-trained general-purpose models, utilizing them as auxiliary tools to enhance OOD detection for domain-specific models is both practical and logical.
One potential limitation of RONIN is its runtime. Diffusion models typically require longer runtimes compared to object detectors. However, this limitation can be mitigated by reducing the number of denoising steps. Since our goal is to detect OOD objects rather than generate high-quality images, a slight compromise in generation quality is acceptable. We examine this trade-off in  section 5 and show that reducing denoising steps significantly decreases runtime with minimal performance degradation.
Moreover, it is worth noting that object detectors are not always deployed in real-time settings. In many real-world scenarios, many real-time imaging often requires offline post processing such as camera-trap biology imaging in-the-wild, where performance and generalizability take precedence over runtime. By demonstrating consistently strong performances across outdoor and indoor scenarios through our experimental design, our framework RONIN effectively showcases its potential in real-time deployment.
Lastly, as generative modeling technologies continue to advance, we expect significant improvements in runtime efficiency, further enhancing the practicality of RONIN. By emphasizing flexibility, generability, and practicality both in our design choice and in our experimental performances, we believe RONIN approach showcases its valuable assets to contribute to the advancement of zero-shot out-of-distribution object detection.
2 Related Work
Out-of-distribution Detection.   The goal of the out-of-distribution (OOD) detection task is to determine whether a given sample belongs to a certain distribution. It has been widely studied at the image level, where a whole image is treated as a sample. Common approaches include leveraging classifier-specific information such as confidence scores (Hendrycks & Gimpel, 2016; Lee et al., 2017; Liang et al., 2018; DeVries & Taylor, 2018; Hsu et al., 2020; Liu et al., 2020; Wei et al., 2022) or learned features (Lee et al., 2018; Denouden et al., 2018; Tack et al., 2020; Sehwag et al., 2021; Xiao et al., 2021; Sun et al., 2022), or directly modeling an image distribution using generative models (Ren et al., 2019; Serrà et al., 2019; Xiao et al., 2020; Schlegl et al., 2017; Zong et al., 2018; Graham et al., 2023; Liu et al., 2023b; Li et al., 2023b). Recent works (Ming et al., 2022; Esmaeilpour et al., 2022; Wang et al., 2023) have also explored using CLIP (Radford et al., 2021) to identify OOD examples in a zero-shot manner, bypassing the need to learn from in-domain data explicitly.
OOD detection can also be extended to the object-level, where objects within an image are treated as individual samples. Most existing research in this setting focuses on training-time interventions. For instance, Du et al. (2022b, a) improve detector features to make them more separable between in-distribution (ID) and OOD data; Wilson et al. (2023) uses adversarial examples to train an MLP for classifying ID and OOD instances. In contrast, our approach RONIN explores the object-level setting through post-hoc interventions instead.
Text-to-Image Generative Models.   Recent advances in generative modeling have made large-scale text-to-image models widely available (Rombach et al., 2021; Ramesh et al., 2022; Saharia et al., 2022). These models exhibit a deep understanding of language and are highly effective at generating or editing high-quality images from diverse prompts, offering a promising approach to data synthesis across various tasks. Furthermore, recent works have shown that these models can also enhance discriminative tasks. For instance, Li et al. (2023a); Jaini et al. (2024) demonstrate that Stable Diffusion (Rombach et al., 2021) can function as an effective zero-shot classifier, achieving accuracy comparable to or surpassing CLIP and various trained discriminative classifiers. There has also been recent progress in applying text-conditioned diffusion models for image-level OOD detection (Du et al., 2023; Gao et al., 2023; Fang et al., 2024). Unlike these works, RONIN focuses on exploring the use of such models for object-level tasks.
3 Problem Formulation
We address the task of object-level out-of-distribution (OOD) detection, which is motivated by the fact that object detectors can recognize unseen objects as seen ones. Specifically, given an object detector trained to detect a pre-defined set of categories (e.g., different kinds of vehicles), we aim to identify the error cases when it wrongly recognizes a novel object (e.g., a wild animal) as one of the pre-defined-categories and detects it overconfidently. We refer to the pre-defined set of categories as the in-distribution (ID) classes, following (Du et al., 2022b); the novel categories as the OOD classes.
Formally, given an image $\bm{x}$ and an object detector $f(.;\ \theta)$ trained for the ID classes $\mathcal{Y}^{in}$, $f(\bm{x};\ \theta)$ outputs a list of bounding boxes $\bm{b}=\{b_{1},\ b_{2},\dots,\ b_{n}\}$ and their associated ID class labels $\bm{\hat{y}}=\{\hat{y}_{1}$, $\hat{y}_{2}$$\dots$, $\hat{y}_{n}$}, where $\hat{y}_{i}\in\mathcal{Y}^{in}$. Object-level OOD detection is then formulated as a binary classification problem, classifying whether the object within $b_{i}$ truly belongs to ID classes. Typically, one would develop a scoring function $\bm{g}$ such that given a bounding box and its predicted label $(b,\ \hat{y})$, $\bm{g}$ gives a higher score $\bm{g}(b,\ \hat{y})$ if $b$ outlines an ID object and a lower score if it outlines an OOD object.
Existing works (Du et al., 2022b, a) proposed a specific training process for the object detector so that its responses (e.g., the feature vector of each bounding box) would better distinguish between ID and OOD classes. However, doing so requires re-training the object detector, assuming access to the original training data and making it infeasible to off-the-shelf object detectors.
Aim.   To address this limitation, we aim to design an OOD detection mechanism that is post-hoc, without the need to modify (e.g., fine-tune) the pre-trained object detector, and zero-shot (Ming et al., 2022), without the need to access the original training data or any ID-class data. This sharply contrasts several prior post-hoc methods that need ID-class data (Ren et al., 2019; Xiao et al., 2020).
Approach.   The emergence of vision-language foundation models (Radford et al., 2021; Li et al., 2022; Liu et al., 2024a) trained on abundant images and free-form text pairs has gradually removed the boundary between the closed-set and open-set settings Li et al. (2023a). For example, while not perfect, CLIP (Radford et al., 2021) can match an image with unbounded concepts. Such a zero-shot capability has been leveraged in prior work to detect OOD samples given a set of ID concepts (Ming et al., 2022; Esmaeilpour et al., 2022; Wang et al., 2023). In this work, we leverage another kind of foundation model, text-to-image generative models (Rombach et al., 2021), capable of generating images given free-form text. While not designed for object detection nor specifically optimized for the ID data, we surmise their built-in, generic capability would facilitate object-level OOD detection.
4 RONIN: Zero-Shot OOD Conditional Inpainting
Figure 2: Overall framework of RONIN with two main components: (i) Class-conditioned Inpainting model and (ii) Vision-language Triplet Measurement. Given an input image with bounding boxes and in-distribution class labels, RONIN masks the objects within each bounding box and conditions the inpainting model to reconstruct them based on their predicted labels. The inpainted outcome is then compared to the original using the vision-language triplet measurement to evaluate semantic and visual alignment, enabling zero-shot OOD detection.
4.1 Theoretical motivation
Our proposed framework, RONIN, is developed based on the assumption that when encountering unseen OOD objects, generative and object detection have different options due to different training objectives. Generally, object detections are trained to separate foreground objects from the background and classify them into one of the pre-defined ID categories—i.e., learning $P(y|x)$. By minimizing this objective, they learn soft decision boundaries between categories, which can overconfidently classify unseen foreground objects into one seen label during inference. In contrast, generally trained generative models aim to model the underlying data distribution $P(x,y)$ per category by capturing their nuanced characteristics. When a sample deviates from the training distribution it learns, it can facilitate identifying this as an outlier. Based on these distinct objectives, we leverage generative models as an auxiliary indicator over object detection outcomes to enable effective zero-shot out-of-distribution detection.
We present the general framework of RONIN in Figure 2. Given an input object with its predicted bounding box and ID label, RONIN utilizes off-the-shelf class-conditioned inpainting models to regenerate the object, conditioning the process by the label to align it with the in-distribution domain. The inpainted outcome is then compared to the original object under a vision-language triplet similarity measurement. High similarity scores reflect strong visual and semantic alignment, suggesting the object is in-distribution, whereas lower scores indicate significant deviation, indicating out-of-distribution. Below, we provide a detailed component of our approach.
4.2 Class-conditional Inpainting
Our key idea behind RONIN is to harness diffusion models for inpainting, refine predicted objects, and bring them closer to the ID domain. Specifically, given a detected bounding box $\bm{b}$ and a predicted ID class $\hat{y}\in\mathcal{Y}^{in}$ by the object detector, we generate a masked region $\bm{m}$ within $\bm{b}$ and use $\hat{y}$ as a condition for reconstruction. Since $\bm{m}$ is always contained within $\bm{b}$, the inpainting model receives sufficient context from both the background and the detected object as hints for reconstructing the original object. This ensures that the inpainted outcomes closely align with the original ID objects, not only in semantic meaning but also in visual appearance, while still distinct in OOD cases.
More specifically, let the semantic background $\bm{s}$ be the entire image region excluding the masked region $\bm{m}$, RONIN takes input of $\bm{m}$ (also known as the inpainting mask), the predicted label $\bm{c}=\hat{y}$, and the semantic background $\bm{s}$ for inpainting. We choose to use pre-trained text-conditioned diffusion models such as DDPM (Ho et al., 2020). The diffusion models starts with a random Gaussian noise over $\bm{m}$ and perform the reverse diffusion process with $T$ steps, guided by $\bm{c}$, to obtain the outcome image $\bm{x}_{inp}={\bm{x}}_{0}$, as in Equation 1, where $\mu_{\theta}$ and $\Sigma_{\theta}$ are priorly learned by pre-training a neural network with parameter $\theta$.
$p_{\theta}(\tilde{\bm{x}}_{t-1}|\bm{x}_{t},\ \bm{c})$ $=\mathcal{N}\big{(}\tilde{\bm{x}}_{t-1};\ \mu_{\theta}(\bm{x}_{t},\ t,\ \bm{c}),\Sigma_{\theta}(\bm{x}_{t},\ t,\ \bm{c})\big{)}$ (1)
$\bm{x}_{t-1}$ $=\bm{\tilde{x}}_{t-1}\odot\bm{m}+\bm{s}$ (2)
While the inpainted objects are synthesized based on $\hat{y}$ over the foreground region $\bm{m}$, the information retained from $\bm{s}$ guides the DDPM to preserve some visual features of the original object, as described in Equation 2. As a result, the inpainted outcome object retains the same visual and semantic characteristics as the original object, despite being presented differently. This enables RONIN to generate objects closely aligned with the originals in ID cases and significantly deviating in OOD cases, as shown in Figure 2. We further analyze the correlation between the bounding box $\bm{b}$ and the inpainted mask $\bm{m}$ under visual and semantic alignment perspectives in Appendix C.
Class-wise Masking and Inpainting.   Intuitively, a natural approach to inpainting is treating each object independently, or “object-wise” inpainting. However, this is highly inefficient when many objects share the same predicted label and denoising condition, making the process redundant. To address this, we propose “class-wise” inpainting, where objects are grouped by the predicted label $\hat{y}$ for a single inpainting process, ensuring the computational cost at most scales with $\mathcal{Y}^{in}$ in the worst case. This significantly reduces computational cost while maintaining performance, as the shared denoising condition ensures effective synthesis. We further evaluate the performance between the two strategies in our ablation study.
4.3 Vision-language Triplet Measurement
Figure 3: Triplet similarity relationships between (i) the original object, (ii) the inpainted outcome, and (iii) the predicted ID label. In-distribution (ID) samples demonstrate strong alignment across all three components, showing high semantic and visual similarity. Conversely, out-of-distribution (OOD) samples exhibit poor alignments, particularly between the original object and the predicted label or inpainted object, effectively differentiating ID from OOD.
Similarity Aligments.   After inpainting, we obtain the original object $\mathbf{x_{ori}}$, its predicted label $\hat{y}\in\mathcal{Y}^{in}$, and the corresponding inpainted object $\mathbf{x_{inp}}$ which is synthesized based on $\hat{y}$. Intuitively, for in-distribution (ID) objects, $\mathbf{x_{inp}}$ tends to closely resemble $\mathbf{x_{ori}}$, while in contrast, out-of-distribution (OOD) objects exhibit lower similarity, indicating divergence from the original given a wrong label.
However, due to the stochastic nature of diffusion inpainting and subtle distinctions in semantic meanings between OOD samples and predicted ID labels, the similarity between $\mathbf{x_{ori}}$ and $\mathbf{x_{inp}}$ can vary greatly. To mitigate this, we introduce a triplet of similarity scores for a more robust ID vs. OOD distinction:
$\textit{similarity}{(ori,\ \hat{y})}$ $=\mathbf{cosine}\left(f_{v}(\mathbf{x_{ori}}),f_{t}(\hat{y})\right)$ (3)
$\textit{similarity}{(inp,\ \hat{y})}$ $=\mathbf{cosine}\left(f_{v}(\mathbf{x_{inp}}),f_{t}(\hat{y})\right)$
$\textit{similarity}{(ori,\ inp)}$ $=\mathbf{cosine}\left(f_{i}(\mathbf{x_{ori}}),f_{i}(\mathbf{x_{inp}})\right)$
Here, $\mathbf{cosine}$ means the cosine similarity between two vectors and $f_{v}$ and $f_{t}$ represent visual and textual embeddings from large vision-language contrastive models (e.g., CLIP (Radford et al., 2021)), while $f_{i}$ is derived from a visual contrastive model (e.g., SimCLRv2 (Chen et al., 2020)). Cosine similarity effectively measures the alignment between these representations. Specifically, $\textit{similarity}{(ori,\ inp)}$ emphasizes visual alignment for ID retention while $\textit{similarity}{(ori,\ \hat{y})}$ supports semantic alignment for OOD recognition. In contrast, $\textit{similarity}{(inp,\ \hat{y})}$ helps normalize variations across different labels. Figure 3 illustrates the relationships between these components for both ID and OOD cases.
Triplet Measurement.   Based on the relationship between the three similarities, we then formulate a triplet OOD score:
$\text{S}_{\text{triplet}}=\frac{\textit{similarity}{(ori,\ \hat{y})}^{\bm{\alpha}}\times\textit{similarity}{(ori,\ inp)}^{\bm{\beta}}}{\textit{similarity}{(inp,\ \hat{y})}}$ (4)
where $\bm{\alpha}$ and $\bm{\beta}$ are hyperparameters controlling the balance between visual-language and visual similarity. A higher $\text{S}_{\text{triplet}}$ suggests a greater likelihood of the object being ID, while lower values indicate OOD cases. In practice, a thresholding method can be applied for effective OOD detection depending on different tasks.
4.4 Refined Prompt for Near-OOD
Identifying near-OOD cases poses a significant challenge for conventional OOD detection methods, as unseen objects often share closely related semantic meanings with certain in-distribution (ID) categories, making separation difficult. However, off-the-shelf generative models inherently possess the ability to distinguish them due to the abundant pre-training data. Given a predefined set of ID labels, we can leverage knowledge graph or taxanomy to retrieve closely related classes or concepts that are OOD and explicitly instruct generative models not to inpaint such concepts as conditioning. For example, if the ID label is “horse”, we can retrieve highly related but OOD classes such as—donkey, zebra, mule, pony, and camel—and explicitly condition the inpainting model to generate a horse that remains distinct from these related categories. This “refined” prompting technique not only guides RONIN to synthesize inpainted outcomes that align better with the ID label but also strengthens its robustness against near-OOD samples.
5 Experiments
Table 1: Object-level OOD detection on four settings. We highlight the best and second best performance. Even with 5 denoising steps, RONIN is able to outperform all the baselines in three out of four settings. Denoising with more steps further improve the performance.
In-Distribution Method MS-COCO OpenImages FPR@95 ()($\downarrow$)( ↓ ) AUROC ()($\uparrow$)( ↑ ) FPR@95 ()($\downarrow$)( ↓ ) AUROC ()($\uparrow$)( ↑ ) PASCAL-VOC ODIN (Liang et al., 2018) 41.65 88.22 55.87 86.46 Energy Score (Liu et al., 2020) 29.48 90.26 24.57 91.24 Mahalanobis (Xiao et al., 2021) 63.30 83.26 45.22 87.11 KNN (Sun et al., 2022) 58.56 85.52 45.00 84.62 VOS (Du et al., 2022b) 48.15 88.75 54.63 83.65 SIREN (Du et al., 2022a) 64.70 78.68 66.69 75.12 MCM (Ming et al., 2022) 62.47 83.15 71.52 81.45 CLIPN (Wang et al., 2023) 43.09 85.45 41.74 89.31 TAG (Liu & Zach, 2024) 61.03 78.35 48.48 87.69 OLE (Ding & Pang, 2024) 54.23 86.13 49.13 88.07 GL-MCM (Miyai et al., 2025) 78.56 78.64 69.35 82.89 RONIN   (5 steps) 29.48 90.39 20.00 92.32 RONIN (10 steps) 29.07 91.10 20.20 92.80 RONIN (15 steps) 25.57 91.63 19.87 92.84 RONIN (20 steps) 25.36 92.31 18.91 93.10 BDD-100k ODIN (Liang et al., 2018) 96.51 55.18 95.56 57.87 Energy Score (Liu et al., 2020) 71.75 74.49 53.33 73.09 Mahalanobis (Xiao et al., 2021) 31.75 90.74 23.33 93.81 KNN (Sun et al., 2022) 35.87 86.58 31.11 92.75 VOS (Du et al., 2022b) 65.45 78.34 59.23 80.42 SIREN (Du et al., 2022a) 42.86 89.37 37.97 91.78 MCM (Ming et al., 2022) 95.56 55.82 92.22 57.05 CLIPN (Wang et al., 2023) 28.49 92.14 44.76 85.78 TAG (Liu & Zach, 2024) 46.43 90.85 55.30 76.67 OLE (Ding & Pang, 2024) 57.14 78.14 40.00 85.36 GL-MCM (Miyai et al., 2025) 96.83 50.82 94.44 58.88 RONIN   (5 steps) 27.94 92.59 26.67 91.23 RONIN (10 steps) 26.67 92.73 25.56 91.33 RONIN (15 steps) 26.03 92.83 24.44 91.65 RONIN (20 steps) 26.03 92.90 23.33 91.90
Figure 4: Visualizations of RONIN’s inpaintings. The ID inpaintings closely resemble the original objects, while the OOD ones appear distinctly different. This allows RONIN to effectively distinguish between ID and OOD objects.
5.1 Experimental Setups
Datasets.   We use Pascal-VOC (Everingham et al., 2010) and Berkerly DeepDrive (BDD-100k) (Yu et al., 2020) as ID data, and MS-COCO (Lin et al., 2014) and OpenImages (Kuznetsova et al., 2020) as OOD data, covering a wide range of indoor and outdoor scenes. We follow the splits from Du et al. (2022a, b), where overlapping categories in the ID and OOD datasets are removed. We select a subset of 200 images from the test set of BDD-100k and 400 each from the test sets of Pascal-VOC, MS-COCO, and OpenImages. Data processing details can be found in Appendix A.
Evaluation Metrics.   We evaluate the OOD detection performance following the two standard OOD metrics: the area under the Receiver Operating Characteristic curve (AUROC) and the false positive rate of OOD objects at 95% true positive rate of in-distribution objects (FPR@95).
Baselines.   We consider three types of baselines: (1) Discriminative zero-shot approaches: MCM (Ming et al., 2022), CLIP-based ODIN (Liang et al., 2018), CLIP-based Energy Score (Liu et al., 2020), CLIPN (Wang et al., 2023), TAG (Liu & Zach, 2024), OLE (Ding & Pang, 2024) and GL-MCM (Miyai et al., 2025). (2) Alternative generative zero-shot approaches: We first generate synthetic data and then apply standard OOD detection methods such as Mahalanobis (Xiao et al., 2021) and KNN (Sun et al., 2022). (3) Training-based approaches: for a complete comparison, we also include two representative training-based object-level methods VOS (Du et al., 2022b) and SIREN (Du et al., 2022a), resulting in a total of eleven baselines. Implementation details can be found in Appendix B.
Implementation Details.    For RONIN, we perform class-wise inpainting with Stable Diffusion 2 Inpainting (Rombach et al., 2021) with 20 steps. Each object is masked with a center mask covering 0.9 of the height and width of its original bounding box. Object-label similarities are determined by the distance between OpenCLIP (Ilharco et al., 2021) features, and object-object similarities by the SimCLRv2 (Chen et al., 2020) features.
For the main experiments in subsection 5.2, we use detected objects obtained from Deformable-DETR (Zhu et al., 2020), as implemented by Du et al. (2022a), to enable direct comparisons with training-based baselines. However, it is important to note that RONIN is compatible with various types of detectors. In subsection 5.3, we demonstrate RONIN ’s performance using detections from the pre-trained open-vocabulary detector GroundingDINO (Liu et al., 2023a), showcasing its flexibility and compatibility. The Triplet Similarity (Equation 4) is computed with $\bm{\alpha}=2$ and $\bm{\beta}=1$.
5.2 Main Experimental Results
Table 1 presents the performance of RONIN and the baselines based on the DeformableDETR detections in four settings. RONIN achieves consistent and competitive performance across all four OOD scenarios, attaining the highest performance in three of them even with just 5 denoising steps. It even outperforms training-based methods VOS and SIREN, demonstrating the effectiveness of using off-the-shelf models to bypass the need for retraining with dedicated in-domain data. Using more denoising steps further improves the performance. In PascalVOC vs. OpenImages, it reduces FPR@95 by 23% over the best-performing baseline. Meanwhile, purely discriminative zero-shot approaches show fluctuating results. Alternative generative approaches combined with standard OOD detection show reasonable performance but still underperform RONIN. This further highlights the advantages of incorporating generative models for object-level OOD detection.
Figure 4 visualizes RONIN’s inpaintings using PascalVOC as the in-distribution dataset. The inpaintings for ID objects closely resemble the original ones, while those for OOD objects appear distinctly different. This allows RONIN to effectively differentiate between the ID and OOD objects.
5.3 Near-OOD Detection
Figure 5: RONIN performances over near-OOD samples with the refined prompting techique. By incorporating closely related species as supplementary exclusions, our approach attains inpaintings that align accurately with the predicted ID labels, thus achieving lower similarity scores for effective near-OOD detection.
To evaluate RONIN ’s ability to handle hard OOD objects, we developed a small near-OOD dataset where ID and OOD objects are both visually and semantically similar. Specifically, we curated 411 samples by selecting ID objects from PascalVOC and OOD objects from MS-COCO and OpenImages, focusing on closely related category ID-OOD pairs such as “horse vs. zebra”, “dog vs. fox”, and “cat vs. raccoon/tiger”. We obtained object predictions using the open-vocabulary object detector GroundingDINO (Liu et al., 2023a), yielding 245 ID objects and 265 OOD objects. As described in subsection 4.4, we apply a “refined” prompting technique that incorporates five pre-anticipated nearest OOD labels as exclusions, with the conditioning prompt for diffusion inpainting illustrated in Figure 5.
Table 2: Near-OOD detection performance. RONIN performs competitively in this setting.
FPR@95 ($\downarrow$) AUROC ($\uparrow$) ODIN (Liang et al., 2018) 19.62 94.87 Energy Score (Liu et al., 2020) 9.81 97.65 MCM (Ming et al., 2022) 31.32 93.65 CLIPN (Wang et al., 2023) 3.77 97.80 TAG (Liu & Zach, 2024) 29.06 94.02 OLE (Ding & Pang, 2024) 3.40 98.33 GL-MCM (Miyai et al., 2025) 44.91 85.66 RONIN (Simple prompting) 7.79 96.86 RONIN (Refined prompting) 5.66 98.03
Table 2 compares RONIN with selected baselines on the near-OOD dataset, showing that our proposed method achieves comparable performance in this challenging scenario, particularly when using the “refined” prompting technique. This improvement is further illustrated in Figure 5, where the additional information in the refined prompt enables RONIN to generate outcomes that align more closely with the predicted ID label, resulting in lower similarity scores and thus better near-OOD detection performance. The distinct qualitative and quantitative improvements between the two prompting techniques strengthen our hypothesis that pre-trained, off-the-shelf generative and multi-modal models already possess the capacity for fine-grained near-OOD detection, given sufficient information.
6 Ablation Studies
Figure 6: Performance-speed tradeoff of RONIN. Smaller image size and fewer denoising steps significantly improve speed with minimal performance loss. At full configuration, RONIN achieves the best performance with reasonable runtime, showcasing its adaptability to efficiency and performance needs. For relative runtime comparison, we incorporate the best-performance baseline, CLIPN, though we do not focus on runtime gaps.
Performance-Speed Trade off.   In Figure 6, we analyze the trade-off between RONIN ’s performance and runtime by varying the number of denoising steps and the image resolutions used during the inpainting process. For relative performance comparison, we include CLIPN (Wang et al., 2023) as the best-performing baseline, though we do not compare absolute runtimes between the two methods. Reducing the number of denoising steps and downscaling image sizes significantly accelerates RONIN, with only minor performance degradation. These results highlight the flexibility of our approach and its potential suitability for low-resource compute environments, with reasonable performance drops.
Table 3: RONIN with object-wise and class-wise inpainting. Class-wise inpainting attain better performance while having faster speed.
FPR@95 ()($\downarrow$)( ↓ ) AUROC ()($\uparrow$)( ↑ ) Object-wise Class-wise Object-wise Class-wise VOC - COCO 26.60 25.36 91.48 92.31 VOC - OpenImages 20.00 18.91 92.68 93.10 BDD - COCO 28.89 26.03 93.74 92.90 BDD - OpenImages 26.67 23.33 92.42 91.90
Class-wise vs. Object-wise Inpainting.    Table 3 compares the performance between “object-wise” inpainting strategy, where each object is masked and inpainted individually, and our default “class-wise” inpainting approach, where objects are processed together based on their labels. By sharing the inpainting process among objects with the same label, class-wise inpainting achieves comparable reconstruction quality while delivering better performance and greater efficiency compared to the object-wise strategy.
Table 4: RONIN performances across different diffusion models. The consistently robust performances across inpainting models, including one-step inpainting, demonstrate RONIN effectiveness without any dependence on specific models or inpainting quality.
Denoising process (steps) COCO / OpenImages Avg runtime/image (seconds) FPR@95 ($\downarrow$) AUROC ($\uparrow$) Kandinsky 2.1 20 28.99 / 23.96 89.92 / 91.62 1.12 Stable Diffusion 1 20 29.90 / 23.70 91.36 / 92.36 1.20 Stable Diffusion 2 (our default) 20 25.84 / 18.91 92.28 / 93.30 0.83 Stable Diffusion 2 5 29.48 / 20.20 90.39 / 92.32 0.19 InstaFlow 1 31.34 / 25.74 88.03 / 90.35 0.03
Reliability of Inpainting.   To assess the stability of RONIN independent of specific diffusion inpainting outcomes, we conducted experiments replacing the default Stable Diffusion 2 with Kandinsky 2.1 and Diffusion Model 1, all using 20 denoising steps. We further included InstaFlow (Liu et al., 2024b) with one-step text-to-image generation, and Stable Diffusion 2 with only 5 steps to assert RONIN with low-step performance. As shown in Table 4, RONIN remains stably robust across different diffusion models and settings. Notably, Stable Diffusion 2 offers the best trade-off between performance and runtime, even with as few as 5 steps. These results suggest that, by subtly exploiting the semantic misalignment between original and synthesized objects for zero-shot OOD detection, our proposed RONIN performs consistently, without relying on any specific diffusion model or high-quality inpainted outcomes. Even when diffusion models produce different inpainting results, our framework remains robust and effective.
Table 5: Ablation studies on the triplet similarity measurement (Equation 4). Results show that all three similarities are essential for providing a stable and effective triplet score (underlined). Additionally, prioritizing the vision-language similarity enhances the performance (bolded).
COCO / OpenImages FPR@95 ($\downarrow$) AUROC ($\uparrow$) without $similarity(inp,\hat{y})$ 44.54 / 37.39 89.77 / 91.42 without $similarity(ori,\hat{y})$ 65.15 / 68.91 69.72 / 78.53 without $similarity(ori,inp)$ 30.93 / 36.96 86.14 / 90.10 triplet similarity, $\bm{\alpha}=\bm{\beta}=1$ 35.46 / 24.78 87.75 / 91.43 triplet similarity, $\bm{\alpha}=1$; $\bm{\beta}=2$ 43.93 / 33.91 85.40 / 89.50 triplet similarity, $\bm{\alpha}=2$; $\bm{\beta}=1$ 25.36 / 18.91 92.31 / 93.10 triplet similarity, $\bm{\alpha}=2$; $\bm{\beta}=2$ 35.46 / 25.43 90.32 / 92.32
Triplet Similarity Measurement.    Table 5 presents ablation studies on the components and hyperparameters of the triplet similarity measurement defined in Equation 4. The upper half of the table (rows 1–4) evaluates the impact of removing individual similarity components, highlighting the necessity of incorporating all components to achieve optimal performance. The lower half (rows 5–7) investigates the balance between semantic alignment ($\textit{similarity}{(ori,\ \hat{y})}$, weighted by $\bm{\alpha}$) and visual alignment ($\textit{similarity}{(ori,\ inp)}$, weighted by $\bm{\beta}$). Given the semantic gap between ID and OOD cases illustrated in Figure 3, emphasizing semantic alignment ($\bm{\alpha}>\bm{\beta}$) leads to optimal performance, resulting in our design choice. For simplicity, we adopt $\bm{\alpha}=2$ and $\bm{\beta}=1$ as our default configuration. However, we hypothesize that the sensitive natures of these weights allow RONIN to adapt flexibly to more challenging scenarios.
7 Discussion
Through consistently strong performance across evaluated experiments and ablation studies, our proposed framework, RONIN, effectively demonstrates its flexibility, generality, and practicality across a wide range of settings and scenarios, from indoor scenes to outdoor or in-the-wild environments. By effectively leveraging the subtle misalignment between generative and discriminative outputs, our method performs robustly without relying on specific diffusion models or inpainting outcomes, while remaining versatile and adaptable to various object detection models, including open-world ones, in a training-free manner. However, RONIN is not without limitations. The computational cost of diffusion models restricts RONIN applicability in real-time settings. Nevertheless, many real-world tasks often require offline post-processing, where accuracy and generalizability are prioritized over runtime, in which our method continues to excel. Moreover, ongoing advancements in generative modeling, such as one-step diffusion models InstaFlow (Liu et al., 2024b), offer promising improvements in efficiency, further enhancing RONIN practicality.
8 Conclusion
Overall, RONIN offers a plug-and-play solution for OOD object detection, delivering strong and consistent performance across scenarios. Without requiring access to ID data or retraining effort, RONIN is highly adaptable and compatible, leveraging general-purpose pre-trained models to enhance domain-specific OOD detection both practically and logically. By subtly leveraging the inconsistency between generative and discriminative models with a triplet similarity that considers both semantic and visual alignments, RONIN excels across various OOD detection tasks and seamlessly integrates with most existing object detectors for flexible, dynamic use. With consistently robust experimental results, RONIN demonstrates its flexibility, generality, and practicality, contributing meaningfully to the advancement of zero-shot out-of-distribution object detection.
Appendix A Data processing
Following Du et al. (2022b, a), we used PascalVOC (Everingham et al., 2010) with 20 categories [^1] PascalVOC contains the following ID labels: Person, Car, Bicycle, Boat, Bus, Motorbike, Train, Aeroplane, Chair, Bottle, Dining Table, Potted Plant, TV Monitor, Couch, Bird, Cat, Cow, Dog, Horse, Sheep.
and BDD100k (Yu et al., 2020) with 10 categories [^2] BDD100k contains the following ID labels: Pedestrian, Rider, Car, Truck, Bus, Train, Motorcycle, Bicycle, Traffic Light, Traffic Sign. We utilized two subsets of MS-COCO (Lin et al., 2014) and OpenImages (Kuznetsova et al., 2020) as the OOD sets, where all images containing ID categories are removed to ensure the disjoint-category requirement. Additionally, for some British labels such as “Aeroplane” or “Couch”, we replaced them with the US counterparts such as “Airplane” or “Sofa” to optimize the usage of CLIP visual-textual representations.
Since our method is zero-shot and does not require training, all the experiments are conducted in subsets taken from the testing set of VOC, BDD, COCO, and OpenImages. To ensure intensive, fairness, and effective evaluations, as well as the balance between ID and OOD objects, we sample a subset of 200 images for BDD-100k testing set and 400 images for PascalVOC, MS-COCO, and OpenImages, as BDD-100k often contain a significantly larger amount of objects per image.
Appendix B Implementation Details for Baselines
For the discriminative zero-shot approaches, MCM (Ming et al., 2022), CLIPN (Wang et al., 2023), TAG (Liu & Zach, 2024), OLE (Ding & Pang, 2024) and GL-MCM (Miyai et al., 2025) are originally developed for image-level OOD detection; we adapt it to the object-level setting by applying it over the cropped detections. ODIN (Liang et al., 2018) and Energy score (Liu et al., 2020) are originally developed for non-zero-shot classifier-based settings; we adapt them to the zero-shot object-level setting by using CLIP to classify the detection crops. For alternative generative approaches, we first generate synthetic objects using Stable Diffusion 2, and then extract features with SimCLRv2 (Chen et al., 2020) and fit OOD detectors using standard methods like Mahalanobis (Xiao et al., 2021) and KNN (Sun et al., 2022), for a fair comparison. For SIREN (Du et al., 2022a) and VOS (Du et al., 2022b), we follow their original settings, using the features from the detectors trained with their proposed schemes.
Appendix C Ablation on Inpainting Masking Ratio
To investigate the correlation between the predicted bounding box and the inpainting mask, we conduct an experiment where the sampled mask covers 25% to 100% of the bounding box. The results, presented in Table 6 and visualized in Figure 7, demonstrate that RONIN is robust across different masking ratios but achieves optimal performance when 75% to 90% of the object is covered. As shown in Figure 7, lower masking ratios result in inpaintings that closely resemble the original objects in both ID and OOD cases, preserving the semantic structure of the originals. Conversely, masking the entire object produces inpaintings that deviate significantly from the originals in both cases, introducing ambiguity in OOD detection. By covering 75%-90% of the object, the inpaintings retain the visual appearance and semantic information of ID objects while generating dissimilar representations for OOD cases, enabling RONIN to effectively distinguish ID and OOD samples.
Table 6: Ablation studies on the masking ratio. RONIN maintains consistent and robust performance across different masking ratios, with the best performance attained when 75%-90% of each object is masked.
Covered masking MS-COCO/OpenImages
ratio $\bm{m}$ FPR@95 ()($\downarrow$)( ↓ ) AUROC ()($\uparrow$)( ↑ )
25% 42.47 / 36.74 88.53 / 90.40
50% 34.23 / 23.91 89.39 / 92.45
75% 30.10 / 21.74 91.43 / 93.41
80% 26.80 / 16.74 91.32 / 93.84
90% 25.57 / 19.57 91.53 / 92.46
100% 28.87 / 21.09 90.61 / 92.61
Figure 7: RONIN inpainting performance with different masking ratios. A small masking ratio preserves much of the original object’s semantic structure, resulting in high similarity for both ID and OOD cases. In contrast, a high masking ratio introduces large visual deviations, leading to lower similarity in both cases. With a balanced masking ratio around 75–90%, RONIN effectively retains the visual and semantic integrity of ID objects while inducing significant deviations in OOD cases, achieving a clear and distinct separation between them.